Sol Summary – October 13th
Author: Alexandre Mangeot
UTC Date: Oct. 13th 2016
Person filling out report: Alexandre MANGEOT, CO
Summary Title: Another routine day on Mars
Mission Status: Operations are near to be nominal. Water consumption needs to be monitored because no resupply is expected for the next 20 days. Furniture are still missing but should arrive tomorrow.
Sol Activity Summary: EVA team has found hypotlith. A sample has been brought back for a first microscope observation. Field scientists are working on future activities and reports.
The current food inventory has been conducted and engineering maintenance and project have been continued.
Anomalies in work: Annalea backpack battery got quickly depleted and had to cut short her EVA.
Weather: Sunny in the morning, cloudy in the afternoon.
Crew Physical Status: No issue.
EVA: Scout for hypolith near the Hab at the road junction.